Assume a corporation has just paid a dividend of $ 4.24 per share. The dividend is expected to grow at a rate of 2.9% per year forever, and the discount rate is 9.3%.

What is the Capital Gains yield of this stock?
Answer #1

Current Dividend = D0 = $ 4.24, Growth Rate of Dividend = 2.9 % and Discount Rate = 9.3 %

Expected Dividend 1 Year from current time = D1 = D0 x (1+g) = 4.24 x 1.029 = $ 4.36296

Expected Dividend 2 Years from current time = D2 = 4.36296 x 1.029 = $ 4.489486

Current Stock Price = P0 = D1 / (Discount Rate - Growth Rate) = 4.36296 / (0.093 - 0.029) = $ 68.17125

Stock Price 1 year from now = P1 = D2 / (Discount Rate - Growth Rate) = 4.489486 / (0.093 - 0.029) = $ 70.14822

Capital Gains Yield = [(70.14822 - 68.17125) / 68.17125] x 100 = 2.9 %